OpenStudy (anonymous):

Write the sum or difference in the standard form a + bi. ( 3 + 3i) - ( -6 + i)

OpenStudy (unklerhaukus):

combine like terms

OpenStudy (anonymous):

( 3 + 3i) - ( -6 + i) = 3+3i +6-i = 3i-i +6+3= 9+2i which is in the form of a+bi where a = 9 and b= 2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Write the expression in standard form. seven divided by quantity three minus fifteen i. What about this?